OpenStudy (anonymous):

HELP ME (: Maya and Sally were the two finalists in a singing competition. The person with the most votes from the audience was chosen as the winner. Sally received 25% of the total votes and lost by 62,500 votes. No person in the audience was allowed to vote more than once. The equation below represents this situation, where y is the total number of votes cast and x is the number of votes Maya received. x - 62,500 = 0.25y If a total of 125,000 people voted, how many votes did Maya receive?

OpenStudy (anonymous):

x - 62,500 = 0.25y x = Maya, y = total, = 125,000 x - 62500 = 0.25(125,000) = 31,250 x - 62500 = 31,250 x = 62500 + 31250 = can you do the rest?

OpenStudy (anonymous):

im kinda confused on this

OpenStudy (anonymous):

just add the last equation x=62500+31250=

OpenStudy (anonymous):

oh ok thanks bro

OpenStudy (anonymous):

u want medal?